# Exact Consistency Under Partial Views: Graph Colorability, Capacity, and Equality in Multi-Location Encodings

### 16 Mar 2026 v3

@@ XIII-B Consistency-Constrained Storage and Interaction
Consistency-constrained storage problems, including multi-version coding and related distributed-storage converses, show that correctness requirements impose unavoidable information costs [13]. Here the object under study is a modifiable encoding architecture, and the main question is the exact-consistency cost of multiple independently writable representations of one latent fact.
XIII-C Computational Realizations
@@
Reflection, metadata systems, and maintenance mechanisms in host platforms provide examples of how the realizability conditions can be instantiated in practice [14, 15]. These examples are auxiliary to the zero-error theory rather than part of its main novelty claim.
XIII-D Formal Verification
@@
The Lean 4 formalization places the work in the tradition of mechanized mathematics and verified theory development [7, 16].
XIV Conclusion
@@
The paper’s main contribution is a structural theory of exact failure under deterministic partial views. Admissible view families induce confusability graphs on latent states; those graphs record which latent alternatives the architecture cannot separate, and they govern the exact recovery law. In the exact full-tuple-space coordinate-view model, this graph class is characterized exactly by upward-closed agreement-set families rather than treated as an arbitrary graph input. In the multi-fact partial-observation extension, exact recovery becomes colorability, exactness on a success set becomes induced-subgraph colorability, and the exact finite weighted-success value is determined by the largest colorable induced subgraph.(L: MFT9, MFT18-24, MFT20, MFT125-131)
@@
Repeated composition preserves that failure structure rather than erasing it: the block-rate sequence is supermultiplicative, its normalized rates converge by Fekete’s lemma to a real asymptotic Shannon capacity equal to the supremum of the finite block rates, and the same capacity is bounded above by complement-chromatic and fixed Lovász-$\vartheta$ upper objects. The one-shot upper invariant is matched with standard orthonormal-representation, primal-PSD, and dual-theta forms.(L: MFT29-30, MFT43-84)
@@
The upper theory is sharp on a structurally characterized subclass. For the original clique-fiber subclass, asymptotic Shannon capacity, the fixed Lovász-$\vartheta$ upper, and the logarithm of the number of fibers coincide. The same equality exports back to the original view-family model through a broader structural route: transitivity of confusability yields the component-cluster collapse, meet-witnessing is a sufficient condition for that transitivity, and fiber coherence is a stronger sufficient condition under which the connected components are exactly the realized transcript fibers.(L: MFT85-91, MFT100-102, MFT136-137)
@@
The same latent-state framework also yields a fact-side affine specialization. The observation-side question asks which latent states are distinguishable from partial views; the fact-side question asks which coordinates determine others across the realized-state family. When the valid states form an affine family, the latter question becomes span membership of coordinate functionals, so the fact indices carry a representable matroid and the minimal determining fact sets are exactly the bases. In this affine regime, the same model carries a graph invariant on states and a matroid invariant on coordinates, and the latter supplies tractable upper bounds for the former. More sharply, the relevant affine rank quantity is exactly the finrank of the image of the restricted coordinate projection, which closes the representation-level link between the matroid language and explicit linear algebra.(L: AFM1-14)
@@
The deterministic finite converse remains the foundation of this theory. Once the observation transcript leaves a nontrivial ambiguity class, the same obstruction can be stated as confusability, counting, conditional-entropy, decoder-output, and finite-gap constraints, together with deterministic data processing and a budgeted finite-error extension. The paper does not claim a new general theorem about arbitrary confusability graphs; rather, it shows that a partial-view encoding model can induce genuinely non-clique failure geometry and reach the classical graph-capacity machinery there. The same model also has a clean boundary theory: rate $1$ is exactly the structural-integrity regime, higher rate makes integrity violations reachable, and realizability plus provenance characterize when that integrity can be certified by the host.
@@
As downstream consequences of that foundation, unit rate is the unique zero-incoherence regime, any higher independent rate makes ambiguity reachable, and the same obstruction yields an $O(1)$ versus $\Omega(n)$ manual-update gap together with a realizability criterion based on causal propagation and provenance observability. Appendix A records representative host-level readings of that boundary-case criterion, and Supplement A adds before/after traces and case-study details showing how the same structural split appears in concrete systems.
This boundary-case classification also has an integrity reading. In Pattern A and Pattern B architectures, the host cannot reliably distinguish a valid derived state from an undetected stale state by its own structural interface. The resulting incoherence is an integrity exposure that leaves stale but plausible states reachable and, in adversarial settings, creates latent attack surface inside the architecture itself.
@@
Computational representation. The graph-theoretic formulation is structurally exponential in the number of facts: for $d$ facts over an alphabet of size $q$, the latent state space has cardinality $q^{d}$, so any explicit confusability graph materialization necessarily starts from an exponentially large vertex set.(L: MFT113) A naive explicit construction ranges over at most $q^{2d}$ ordered state pairs, formalized in the artifact through the corresponding product-cardinality and pair-count bounds.(L: MFT116-117) At the same time, the graph need not be materialized: adjacency is already packaged by an implicit oracle on state pairs, formalized as a Boolean confusability test equivalent to the confusability relation itself, and also by an equivalent agreement-set oracle whose direct scan cost is bounded linearly in the total view size.(L: MFT114-115, MFT132-135) Under the explicit coordinate-view representation used throughout the paper, that one-shot oracle is therefore polynomial in its direct input size, computable by scanning the coordinates and the admissible view family in $O(d+\sum_{\ell}|V_{\ell}|)$ time, hence $O(Ld)$. In the affine specialization, the upper bounds are likewise polynomial-time computable once the direction space is presented explicitly by a basis or generator matrix, because the formalized rank quantity is exactly the finrank of the image of the restricted coordinate projection, so each bound reduces to Gaussian elimination on that explicit map.(L: AFM10-11) The exponential barrier therefore lies in full graph materialization and global capacity computation, not in the local adjacency test or in the affine upper-bound surrogate.
Limitations and future work. The present paper develops only the finite deterministic architectural arc. For the exact tuple-space coordinate-view model, the full labeled-tuple-space graph class is now characterized exactly by monotone agreement-set families; what remains open is the corresponding classification up to unlabeled graph isomorphism, the restricted-state and stochastic-support variants, and the sharpness of the upper theory on those broader subclasses.
XIV-A Artifacts
@@
The Lean 4 formalization is included as supplementary material. Appendix A records representative realizability readings, and Supplement A contains the companion case-study material.
@@ Appendix A System Classification
This appendix provides a consolidated classification of representative systems according to the architectural patterns identified by Corollary XII.10. The pattern labels used in the table are: Pattern A (Blind Update), meaning missing causal propagation; Pattern B (Amnesiac Derivation), meaning missing provenance observability; and Pattern C (Coherent Kernel), meaning both conditions are present and verifiable rate $1$ is realizable.
Host family Host / Pattern Class Prop. Prov. Rate 1 Mechanism
@@
Prog. languages Python [17, 18] C yes yes yes For the structural facts considered here, class-definition hooks and runtime hierarchy introspection are host-native.
@@
Databases Engine-maintained materialized view [19, 20] C yes yes yes The engine maintains the derived relation and exposes catalog metadata for it.
@@
Prog. languages Rust [21] B partial no no Compile-time macro generation exists, but the compiled runtime artifact does not retain derivation provenance.
@@
Prog. languages TypeScript / JavaScript [22, 23] B partial no no Decorators can attach metadata, but type erasure and missing subclass provenance block exact verification in the host runtime.
@@
Prog. languages Java [24] A/B no no no Under the runtime-only host interpretation used here, annotations and related derivation tooling are external to the JVM runtime.
@@
Prog. languages Go [25] A/B no no no Reflection exists, but there are no definition-time hooks or host-native implementer registries.
@@
Databases External ETL copy / duplicate summary table A/B no no no Synchronization is externalized, and provenance is no longer intrinsic to the host database.
@@
Deps. npm [26, 27, 28] A no yes no Lockfile and query commands expose provenance, but manifest edits require explicit npm operations before the derived state is updated.
@@
Deps. Cargo [29, 30, 31] A no yes no Resolved-dependency provenance is exposed, but Cargo.lock remains a distinct derived artifact synchronized through Cargo commands.
@@
Deps. Poetry [32] A no yes no The resolved graph is queryable, but lock regeneration is an explicit step rather than part of the source edit itself.
@@
Deps. pnpm [33, 34, 35] A no yes no Provenance is queryable, but the lockfile remains separately maintained rather than automatically updated with source edits.

### 31 Mar 2026 v4

@@ XII-B Consistency-Constrained Storage and Interaction
Consistency-constrained storage problems, including multi-version coding and related distributed-storage converses, show that correctness requirements impose unavoidable information costs [34]. The consistency literature establishes fundamental tradeoffs between consistency, availability, and partition tolerance [35, 36, 37], while correctness models such as linearizability specify formal guarantees for concurrent operations [38]. Here the object under study is a modifiable encoding architecture, and the main question is the exact-consistency cost of multiple independently writable representations of one latent fact.
XII-C Computational Realizations
@@
Reflection, metadata systems, and maintenance mechanisms in host platforms provide examples of how the boundary-case realizability conditions can be instantiated in practice [39, 40]. These examples are secondary to the paper’s main zero-error and graph-capacity contribution.
XII-D Formal Verification
@@
The Lean 4 formalization places the work in the tradition of mechanized mathematics and verified theory development [9, 10, 41].
XIII Conclusion
@@
The paper’s main contribution is a zero-error theory for deterministic partial views. Admissible view families induce confusability graphs on latent states; those graphs record exactly which latent alternatives the architecture cannot separate, and they govern the exact recovery law. In the exact full-tuple-space coordinate-view model, the realizable confusability relations are characterized exactly by upward-closed agreement-set families. In the multi-fact extension, exact recovery becomes graph colorability, exactness on a success set becomes induced-subgraph colorability, and the exact finite weighted-success value is determined by the largest colorable induced subgraph.757575: MFT9, MFT18-24, MFT20, MFT125-131
@@
Repeated composition preserves that failure structure rather than erasing it: the block-rate sequence is supermultiplicative, its normalized rates converge by Fekete’s lemma to a real asymptotic Shannon capacity equal to the supremum of the finite block rates, and the same capacity is bounded above by complement-chromatic and fixed Lovász-$\vartheta$ upper objects. The one-shot upper invariant is matched with standard orthonormal-representation, primal-PSD, and dual-theta forms.767676: MFT29-30, MFT43-84
@@
The upper theory is sharp on a structurally characterized subclass. For the original clique-fiber subclass, asymptotic Shannon capacity, the fixed Lovász-$\vartheta$ upper, and the logarithm of the number of fibers coincide. The same equality exports back to the original view-family model through a broader structural route: transitivity of confusability yields the component-cluster collapse, meet-witnessing is a sufficient condition for that transitivity, and fiber coherence is a stronger sufficient condition under which the connected components are exactly the realized transcript fibers.777777: MFT85-91, MFT100-102, MFT136-137
@@
The same latent-state framework also yields a fact-side affine specialization. The observation-side question asks which latent states are distinguishable from partial views; the fact-side question asks which coordinates determine others across the realized-state family. When the valid states form an affine family, the latter question becomes span membership of coordinate functionals, so the fact indices carry a representable matroid and the minimal determining fact sets are exactly the bases. In this affine regime, the same model carries a graph invariant on states and a matroid invariant on coordinates, and the latter supplies tractable upper bounds for the former. More sharply, the relevant affine rank quantity is exactly the finrank of the image of the restricted coordinate projection, which closes the representation-level link between the matroid language and explicit linear algebra.787878: AFM1-14
@@
The deterministic finite converse remains the foundation of this theory. Once the observation transcript leaves a nontrivial ambiguity class, the same obstruction can be stated as confusability, counting, conditional-entropy, decoder-output, and finite-gap constraints, together with deterministic data processing and a budgeted finite-error extension. The paper does not claim a new theorem about arbitrary confusability graphs; rather, it identifies a deterministic partial-view model that induces a fully characterized nontrivial graph class and carries the classical zero-error graph-capacity machinery with it.
@@
As boundary consequences, unit rate is the unique zero-incoherence regime, any higher independent rate makes ambiguity reachable, and the same obstruction yields an $O(1)$ versus $\Omega(n)$ manual-update gap together with a realizability criterion based on causal propagation and provenance observability. Appendix A records representative host-level readings of that boundary-case criterion together with companion case-study details.
@@
Computational representation. The graph-theoretic formulation is structurally exponential in the number of facts: for $d$ facts over an alphabet of size $q$, the latent state space has cardinality $q^{d}$, so any explicit confusability graph materialization necessarily starts from an exponentially large vertex set.797979: MFT113 A naive explicit construction ranges over at most $q^{2d}$ ordered state pairs, formalized in the artifact through the corresponding product-cardinality and pair-count bounds.808080: MFT116-117 At the same time, the graph need not be materialized: adjacency is already packaged by an implicit oracle on state pairs, formalized as a Boolean confusability test equivalent to the confusability relation itself, and also by an equivalent agreement-set oracle whose direct scan cost is bounded linearly in the total view size.818181: MFT114-115, MFT132-135 Under the explicit coordinate-view representation used throughout the paper, that one-shot oracle is therefore polynomial in its direct input size, computable by scanning the coordinates and the admissible view family in $O(d+\sum_{\ell}|V_{\ell}|)$ time, hence $O(Ld)$. In the affine specialization, the upper bounds are likewise polynomial-time computable once the direction space is presented explicitly by a basis or generator matrix, because the formalized rank quantity is exactly the finrank of the image of the restricted coordinate projection, so each bound reduces to Gaussian elimination on that explicit map.828282: AFM10-11 The exponential barrier therefore lies in full graph materialization and global capacity computation, not in the local adjacency test or in the affine upper-bound surrogate.
Limitations and future work. The present paper develops only the finite deterministic architectural arc. For the exact tuple-space coordinate-view model, the full labeled-tuple-space graph class is now characterized exactly by monotone agreement-set families. Several directions remain open. First, the classification up to unlabeled graph isomorphism is incomplete: two distinct upward-closed families may yield isomorphic confusability graphs, and a complete invariant-theoretic characterization of the model-generated graph class is not yet available. Second, the restricted-state case, where the realized latent tuples form a proper subset of the full product space, admits a richer confusability structure whose graph-theoretic properties are not fully resolved. Third, the stochastic support-overlap extension, in which each location may emit any observation from a finite support set and two states are confusable when some location has overlapping support, is formalized in the Lean artifact but not developed textually; its graph-theoretic consequences and the sharpness of the Lovász-$\vartheta$ upper bound on the resulting broader graph class remain open. Finally, the sharpness of the capacity upper bounds on the deterministic model’s non-clique subclass is not established: the binary square achieves capacity-$\vartheta$ equality by coincidence rather than by the transitivity mechanism of Section VII, and the general conditions under which the upper theory is tight for the model-generated class are not known.
XIII-A Artifacts
@@
The Lean 4 formalization is included as supplementary material. Appendix A records representative realizability readings and companion case-study material.
@@ Appendix A System Classification
This appendix applies the realizability criterion of Corollary XI.10 to representative channel witnesses. Table A classifies systems by their structural side-information capabilities and zero-delay state synchronization properties. Pattern A: missing causal propagation. Pattern B: missing provenance observability. Pattern C: both conditions present (verifiable rate $1$).
Channel Instance Class Prop. Prov. Rate 1 Mechanism
@@
Prog. languages Python [42, 43] C yes yes yes Class-definition hooks and runtime hierarchy introspection are host-native.838383: PYH1, PYI1
@@
Databases Engine-maintained materialized view [44, 45] C yes yes yes Engine maintains the derived relation and exposes catalog metadata.848484: DBV1-3
@@
Prog. languages Rust [46] B partial no no Compile-time macro generation exists, but the compiled runtime artifact does not retain derivation provenance.858585: RUH1, RUI1
@@
Prog. languages TypeScript / JavaScript [47, 48] B partial no no Decorators can attach metadata, but type erasure and missing subclass provenance block exact verification.868686: TSE1, TSN1
@@
Prog. languages Java [49] A/B no no no Under the runtime-only host interpretation used here, annotations are external to the JVM runtime.878787: JVA1
@@
Prog. languages Go [50] A/B no no no Reflection exists, but there are no definition-time hooks or host-native implementer registries.888888: GOL1
@@
Databases External ETL copy / duplicate summary table A/B no no no Synchronization is externalized, and provenance is no longer intrinsic to the host database.898989: DBE1-3
@@
Deps. npm [51, 52, 53] A no yes no Lockfile (auxiliary tag) and query commands expose provenance, but manifest edits require explicit out-of-band operations.909090: NPM1-3
@@
Deps. Cargo [54, 55, 56] A no yes no Resolved-dependency provenance is exposed, but the auxiliary tag (Cargo.lock) remains a distinct derived artifact.919191: CRG1-3
@@
Deps. Poetry [57] A no yes no The resolved graph is queryable, but lock regeneration is an explicit step rather than part of the source edit itself.929292: POT1-3
@@
Deps. pnpm [58, 59, 60] A no yes no Provenance is queryable, but the lockfile remains separately maintained rather than automatically updated.939393: PNM1-3